• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Punjabi

    Log entry time 16:13:04 on October 23,1999

    Entry number 28924

    keyword=shift summary

    8:00 to 830 finished elastic ep run for exp. 94-012
    8:30 to 10:20 did Moller
    10:20 started laser change in the injector to get more current (about 100 micro amps) for exp. 94-004.
    During the time Moller was beging done changed angle and momentum in both spectrometers for exp 94-004 first point.
    Hadron angle = 58.733 and electron angle = 15.995.
    Hadron momentum = 0.890 GeV/s and Electron momentum = 2.760 GeV/c.
    Ready to start the experiment when we get the beam.
    13:45 started doing hHARP scan
    14:45 did several harp scans (harp run 246 to 251). best MCC can do is horizomntal beam width at BPM 3B is 0.55 mm. The experiment 94-004 wants about 0.2 mm.
    MCC calling optics expert.
    14:55 now getting ready to do BCM calibration
    15:00 BCM calibration canceled.
    16:00 still doing harp scan. Horizontal beam width now 0.11 mm (ok) but vertical is about 0.4 and so MCC is still working to improve it.